OpenStudy (tester97):

The 5th term in a geometric sequence is 140. The 7th term is 35. What are the possible values of the 6th term of the sequence?

OpenStudy (ankit042):

ar^4=140 ar^6=35 r^2=35/140

hartnn (hartnn):

Alternate method : \(a_6\) will be the geometric mean between a5 and a7 so, you can simply do \(\large a_6 ^2 = a_5\times a_7\) take square root and don't forget to consider the negative root too

OpenStudy (tester97):

So it would be \[\pm70\]

hartnn (hartnn):

correct! :)

OpenStudy (tester97):

OMG thanks hartnn :)

hartnn (hartnn):

did i ever give you this ? http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0 might be helpful to you now...

OpenStudy (tester97):

MY goodness its like a candy shop o.o thanks ^_^